OpenStudy (anonymous):

help plz evaluate 5P2 a 20 b 30 c 60 d 120 thanks

OpenStudy (anonymous):

what am i evaluate

OpenStudy (anonymous):

idk are u

OpenStudy (anonymous):

i evaluate 5P2 in got this 5p2 =5*p2 =5*(p*p) =5p2

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

your welcome in if that help this should Evaluate for p=5 5(52) 5(52) =125